OpenStudy (anonymous):

If f(x) = 1 – x, which value is equivalent to |f(i)|? options are: A.0 B.1 C. √2 D. √-1

OpenStudy (freckles):

\[|a+bi|=\sqrt{a^2+b^2}\]

OpenStudy (chrisdbest):

B should be the closest

OpenStudy (freckles):

it actually isn't B

OpenStudy (chrisdbest):

Bcuz i = sqrt -1, and the absolut value of a negative number is a positive number

OpenStudy (freckles):

\[f(i)=1-i \\ |f(i)|=|1-i|=?\]

OpenStudy (anonymous):

sorry guys but im still in the blue here ahahah

OpenStudy (chrisdbest):

Oh so its A!!!

OpenStudy (freckles):

\[|1-i|=\sqrt{1^2+1^2}=?\]

OpenStudy (chrisdbest):

Cuz 1 - 1 = 0!

OpenStudy (chrisdbest):

Bam!! Medal me please

OpenStudy (freckles):

or if you wanted to look at it as \[\sqrt{1^2+(-1)^2} \text{ instead of } \sqrt{1^2+1^2}\]

OpenStudy (anonymous):

ayyyy lmao

OpenStudy (freckles):

all I'm asking you to do is do the 1+1 underneath the radical

OpenStudy (freckles):

\[|a+bi|=\sqrt{a^2+b^2} \ \\ \text{ and you have } \\ |1+-1i|=\sqrt{1^2+(-1)^2}=\sqrt{1^2+1^2}=?\] can you finish this ?

OpenStudy (anonymous):

OOOOOOOO THANKS MAN LMAO I DONT KNOW HOW I DIDNT SEE THAT ahha its √2

OpenStudy (freckles):

yep :)

OpenStudy (chrisdbest):

Dang I didn't even see that answer choice. I'm Dyslexic, I thought it said \[\sqrt{21}\]